1. It's ok to use an 29 ga., 3/8" insulin syringe to administer other drugs subcutaneous besides insulin. TRUE OR FALSE?

2. Before withdrawing insulin from the bottle or using a prefilled device, the container should be shaken vigorously to make sure the suspension is evenly distributed.

TRUE OR FALSE?

3. Levothyroxine (Synthroid) 150mcg PO per day is ordered for a patient with hypothyroidism. It is supplied as 0.075 mg tablets. How many tablets should the patient take daily?________tablets

4. Dan must drink 2250 cc of water each day. How many liters of water will he drink?
Round your answer to the nearest tenth liter.

5. A child who weighs 22 pounds is prescribed levothyroxine (Levothroid) 10 mcg/kg daily. How many mcg will the child receive each day? _____ mcg
Answer #1

1. It's ok to use a 29 ga., 3/8" insulin syringe to administer other drugs subcutaneous besides insulin.

False.

Usually, an insulin syringe is not recommended to administer any other drugs.

2. Before withdrawing insulin from the bottle or using a prefilled device, the container should be shaken vigorously to make sure the suspension is evenly distributed.

False

Rationale:

Insulin or prefilled device should not be shaken vigorously before withdrawing because it damages the molecules. The correct way is to keep the insulin between two palms and gently load to resuspend the solution.

3. Levothyroxine (Synthroid) 150mcg PO per day is ordered for a patient with hypothyroidism. It is supplied as 0.075 mg tablets. How many tablets should the patient take daily? 2 tablets.

Ordered dose = 150mcg

Dose on hand = 0.075mg

1mg = 1000mcg

Therefore, 0.075mg = 0.075 X 1000

= 75mcg

1tablets contains 75mcg.

To get 150mcg = 150 / 75

To get = 2 tablets

4. Dan must drink 2250 cc of water each day. How many liters of water will he drink? 2.2L

1litre = 1000cc

2250cc = 2250 / 1000

=2.25 Litres.

Dan drink of 2.2litres.(roundest to the nearest tenth)

5. A child who weighs 22 pounds is prescribed levothyroxine (Levothroid) 10 mcg/kg daily. How many mcg will the child receive each day? 100mcg

1 kg = 2.2lbs

22 pounds = 22/ 1.2

= 10kg

Ordered dose of Levothyroxine = 10mcg /kg

Child weigh's = 10kg

Dose = 10 X 10

Dose = 100mcg day

The child receives 100mcg of Levothyroxine per day.